An unruly demonstrator lifts a stone of mass 200 g from the ground and throws it at his opponent. At the time of projection, the stone is 150 cm above the ground and has a speed of 3.00 m/s. Calculate the work done by the demonstrator during the process. If it takes one second for the demonstrator to lift the stone and throw, what horsepower does he use?

Grade:upto college level

1 Answers

Aditi Chauhan
9 years ago
Sol. m = 200g = 0.2kg, h = 150cm = 1.5m, v = 3m/sec, t = 1 sec Total work done = ½ mv^2 + mgh = (1/2) × (0.2) ×9 + (0.2) × (9.8) × (1.5) = 3.84 J h.p. used = 3.84/746 = 5.14 * 10^-3